OpenStudy (anonymous):

Anu, Manu & Kanu invest Rs.30,000, Rs.20,000 & Rs.15,000 in cloth business. Anu & Manu get respectively 15% & 5% of total profit. The remaining profit is divided among the three friends in the ratio of their investments. If Anu gets Rs.800 more than Manu, find Anu’s share in the profit.

OpenStudy (shrutipande9):

@hartnn would be able to help u better....

OpenStudy (anonymous):

An amount of money grows up to Rs 11520 in 2 years and up to Rs 13824 in 3 years on compound interest. What is the sum?

OpenStudy (anonymous):

i found r=20% then i have to use compound interest formula?

OpenStudy (shrutipande9):

@divu.mkr

OpenStudy (anonymous):

11520=p(1+20/100)2

OpenStudy (anonymous):

8000

OpenStudy (shrutipande9):

srry srry i was wrong... A formula for calculating annual compound interest is A = P \left(1 + \frac{r}{n}\right)^{nt} where A = value after t periods P = principal amount (initial investment) r = annual nominal interest rate (not reflecting the compounding) n = number of times the interest is compounded per year t = number of years the money is borrowed for

OpenStudy (anonymous):

u r right @shrutipande9

OpenStudy (shrutipande9):

\[A=P+(1+\frac{ r }{ n })^{nt}\]

OpenStudy (anonymous):

i have one more question on simple interest

OpenStudy (shrutipande9):

aahh...dat i can do...

OpenStudy (anonymous):

A certain amount of money becomes five times at the simple interest rate of 7% per annum. At what rate per cent will it become 9 times?

OpenStudy (shrutipande9):

so let d principle be x. so amount would be 5x. so intrest would be A-P=4x. now use d formula I=PNR/100. SIMILARLY FOR THE 9TIMES. FIND N and then equate the 2 N's to find x

OpenStudy (anonymous):

4x=x*n*7/100 400=7xn n=400/7

OpenStudy (anonymous):

14% =r got it :-)

OpenStudy (anonymous):

for this partnership question

OpenStudy (shrutipande9):

Cool:D